### v1

# ThinkQE: Query Expansion via an Evolving Thinking Process

Abstract
Effective query expansion for web search benefits from promoting both exploration and result diversity to capture multiple interpretations and facets of a query. While recent LLM-based methods have improved retrieval performance and demonstrate strong domain generalization without additional training, they often generate narrowly focused expansions that overlook these desiderata. We propose ThinkQE, a test-time query expansion framework addressing this limitation through two key components: a thinking-based expansion process that encourages deeper and comprehensive semantic exploration, and a corpus-interaction strategy that iteratively refines expansions using retrieval feedback from the corpus. Experiments on diverse web search benchmarks (DL19, DL20, and BRIGHT) show ThinkQE consistently outperforms prior approaches, including training-intensive dense retrievers and rerankers. [^1] Our code is publicly available at https://github.com/Yibin-Lei/Think_QE.
ThinkQE: Query Expansion via an Evolving Thinking Process
Yibin Lei1, Tao Shen2, Andrew Yates3 1University of Amsterdam  2University of Technology Sydney 3Johns Hopkins University, HLTCOE y.lei@uva.nl, tao.shen@uts.edu.au, andrew.yates@jhu.edu
1 Introduction
Query expansion (QE) is a common practice in web search scenarios Robertson (1990); Qiu and Frei (1993), particularly for first-stage retrievers such as BM25 (Robertson et al., 1995). Effective expansion involves not only reinforcing the core intent of the query but also introducing terms that capture different facets or interpretations of the information need. This multifaceted coverage helps capture a broader semantic context, enabling the retrieval of a more comprehensive set of relevant documents. Prior studies have shown that such broad-coverage expansion strategies lead to substantial improvements in retrieval quality Bouchoucha et al. (2013).
Recent advances in large language models (LLMs) have led to strong performance in query expansion Gao et al. (2022); Wang et al. (2023); Jagerman et al. (2023); Mackie et al. (2023); Shen et al. (2024); Lei et al. (2024), particularly due to their ability to rapidly adapt to new domains without requiring additional training. However, existing LLM-based methods often pay limited attention to exploration and result diversity. As illustrated in Table 1, we observe that current approaches – such as HyDE – tend to generate overly confident expansions that focus narrowly on a single interpretation of the input query. This behavior can be attributed to the model’s reliance on its internal knowledge and high-probability completions Yona et al. (2024); Ohi et al. (2024); Sun et al. (2025), which may suppress alternative formulations or less common aspects of the query. This lack of breadth limits the retrieval of documents reflecting alternative scenarios or requiring more nuanced reasoning.
Query: Who is robert gray
Expansion w/o. Thinking:
Robert Gray is best known as the American captain who discovered the Columbia River in 1792. He named the river after his ship, the Columbia Rediviva, and explored it up to Grays Bay. His discovery was later documented by Lieutenant William Broughton during the Vancouver expedition.
ThinkQE:
Robert Gray is best known as Captain Robert Gray, an American explorer who played a significant role in the exploration of the Pacific Northwest. In 1792, he captained the ship Columbia Rediviva and became the first American to navigate the Columbia River, which he named after his vessel. On May 11, 1792, he entered the mouth of the river and explored approximately 20 miles upstream as far as Grays Bay, which was later named in his honor by Lieutenant William Broughton of the Vancouver expedition. This expedition contributed to the mapping and understanding of the region, highlighting Gray’s importance in early American exploration.
Table 1: Examples comparing a standard expansion with ThinkQE, our proposed query expansion method with thinking-augmentation. ThinkQE encourages deeper reasoning and multifaceted contextualization.
To address these limitations, we propose ThinkQE, a new framework that improves exploration and result diversity along two complementary dimensions. First, we introduce a thinking-based expansion process, where the model explicitly accumulates intermediate thoughts and hypotheses before producing final expansions. This encourages the emergence of new and more exploratory terms that can help retrieve documents beyond the initial query scope. Second, inspired by pseudo-relevance feedback Rocchio Jr (1971), we propose an interactive expansion strategy, where query expansions are progressively refined using feedback from the documents retrieved at each stage. This dynamic interaction with the corpus allows the query to evolve in a context-aware manner, adapting to newly retrieved evidence.
By combining both components, we develop ThinkQE, a test-time query expansion method that achieves strong performance on natural language web search benchmarks, including DL19, DL20, and the StackExchange domain of the BRIGHT benchmark [^2] We omit math and coding subsets, as ThinkQE relies on natural language expansions, which may not be well-suited for symbolic or structured domains.. Remarkably, ThinkQE requires no additional training, yet surpasses recent training-intensive reranking methods, including those based on reinforcement learning and distillation from DeepSeek-R1. Our analysis reveals that: (1) explicitly modeling a thinking process enhances expansion quality, and (2) iteratively refining queries with evolving retrieval feedback is more effective than generating static expansions, even under the same compute budget.
2 Method
We introduce ThinkQE, a query expansion framework that tightly integrates LLM-based thinking process with evolving corpus interaction. The method is designed to enable exploration of the query space through thinking processes and evolving refinement based on retrieval feedback from the corpus. The overall process proceeds in multiple rounds. At each round, an LLM performs thinking-augmented expansion based on the original query and newly retrieved documents from the corpus, which in turn informs subsequent retrieval and expansion steps. This section describes each component of the method in detail.
2.1 Retrieving Initial Evidence from Corpus
Let $q_{0}$ denote the original user query. To ground the expansion process in corpus evidence, we begin by retrieving an initial set of documents from the corpus $\mathcal{C}$ using a first-stage lexical retriever. In our implementation, we employ BM25. Specifically, we retrieve the top-$K$ documents: $\mathcal{D}_{0}=\text{TopK}(\text{BM25}(q_{0},\mathcal{C})).$
Here, $\mathcal{D}_{0}$ denotes the ranked list of top-$K$ documents retrieved for $q_{0}$, ordered by their BM25 relevance scores. This list serves as the initial feedback signal for expansion, providing retrieval-grounded context to the LLM in the first expansion step.
2.2 Expansion via Thinking Process
To produce an initial expansion, we use an R1-distilled LLM trained to generate a thinking chain before answering. Given the original query $q_{0}$ and top-$K$ retrieved documents $\mathcal{D}_{0}$, the model follows a two-phase process:
1. Thinking Phase: The model reflects on $q_{0}$ and $\mathcal{D}_{0}$ to identify latent concepts, resolve ambiguities, and surface alternative interpretations or missing aspects of the information need.
2. Expansion Phase: Based on the thinking output, the model generates a query expansion segment $e_{1}$ that builds upon the original query by introducing additional relevant terms and concepts.
Leveraging the R1-distilled model’s natural separation of thought and answer allows us to implement the reasoning-expansion workflow without additional scaffolding or prompt engineering. The prompt shown in Table 2 guides the model to generate expansions by thinking over the input query and the top-retrieved documents.
ThinkQE Prompt
Given a question "{$q$}" and its possible answering passages (most of these passages are wrong) enumerated as:
1. $\{d_{1}\}$; 2. $\{d_{2}\}$; 3. $\{d_{3}\}$
please write a correct answering passage. Use your own knowledge, not just the example passages!
[…]
2.3 Evolution via Corpus Feedback
We propose to iterate the above thinking-based expansion. At each round $t=1,\dots,T$, the method performs the following steps:
1. Retrieval: The current query $q_{t}$ is used to retrieve a ranked list of documents from the corpus: $\mathcal{R}_{t}=\text{BM25}(q_{t},\mathcal{C}).$
2. Redundancy Filtering: To promote diversity and avoid repetition, we exclude documents that (a) appear in the blacklist $\mathcal{B}_{t}$, or (b) were among the top-$K$ results in the previous round $\mathcal{D}_{t-1}$. We then select the top-$K$ documents from the remaining candidates: $\mathcal{D}_{t}^{\text{new}}=\text{TopK}(\mathcal{R}_{t}\setminus(\mathcal{B}_{t}\cup\mathcal{D}_{t-1})).$ The blacklist is updated to include all documents that were filtered out in this round.
3. Expansion via Thinking: The LLM is prompted with the original query $q_{0}$ and the filtered document set $\mathcal{D}_{t}^{\text{new}}$ to generate the next expansion $e_{t+1}$, using the same two-phase expansion process described in Section 2.2.
4. Query Update: The query is iteratively updated by concatenating the new expansion: $q_{t+1}=q_{t}\oplus e_{t+1}$.
This loop can be repeated for any number of rounds $T$, depending on resource constraints or desired depth.
[…]
Remark.
Our method introduces two key innovations: (1) the use of an explicit, LLM-guided thinking process to encourage deeper exploration during expansion, and (2) an evolving loop that dynamically refines the query based on retrieval feedback. Within this evolving process, we design two essential components – redundancy filtering and expansion accumulation – both of which play a critical role in the effectiveness of ThinkQE, as demonstrated in our results in Section 4.3.
3 Experiments
3.1 Setup
Datasets.
We evaluate ThinkQE on two categories of natural language web search datasets: (1) Factoid-style retrieval: TREC DL19 (Craswell et al., 2020) and DL20 (Craswell et al., 2021), widely used benchmarks based on the MS MARCO document collections (Bajaj et al., 2016); and (2) Reasoning-oriented datasets: The StackExchange domain of the BRIGHT benchmark (Su et al., 2025), covering seven diverse sub-domains: Biology (Bio.), Earth Science (Earth.), Economics (Econ.), Psychology (Psy.), Robotics (Rob.), Stack Overflow (Stack.), and Sustainable Living (Sus.). On the BRIGHT benchmark, we omit math and coding datasets to focus on only the StackExchange subsets, as ThinkQE relies on language-model-based natural language expansions, which may not be well-suited for symbolic or structured domains such as code or math.
Implementation.
We use the QWEN-R1-Distill-14B model (DeepSeek-AI, 2025) to generate thinking-based query expansions, sampling outputs with a temperature of 0.7. The BM25 retrieval is performed using Pyserini Lin et al. (2021) with default hyperparameters. At each round, ThinkQE uses the top-5 retrieved documents (truncated to 128 tokens for DL benchmarks and 512 tokens for BRIGHT) to prompt the LLM, and samples 2 candidate expansions to enhance diversity. We set the total number of interaction rounds to 3, for a balance between efficiency and effectiveness.
DL19 DL20
mAP ndcg@10 R@1k mAP ndcg@10 R@1k
BM25 30.1 50.6 75.0 28.6 48.0 78.6
Supervised Fine-Tuned Dense retrievers
DPR 36.5 62.2 76.9 41.8 65.3 81.4
ANCE 37.1 64.5 75.5 40.8 64.6 77.6
[…]
R1-Distilled Rerankers on BM25 Top-20 Docs
Rank1-32B - 64.9 - - 61.2 -
Rank-K-32B - 66.2 - - 64.3 -
Zero-shot Query expansions with BM25
HyDE 41.8 61.3 88.0 38.2 57.9 84.4
Query2doc - 66.2 - - 62.9 -
MILL - 63.8 85.9 - 61.8 85.3
LameR 42.8 64.9 84.2 - - -
ThinkQE-14B 45.9 68.8 89.3 43.9 64.7 87.8
[…]
Training Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
BM25 Zero-shot 18.2 27.9 16.4 13.4 10.9 16.3 16.1 17.0
BM25 + GPT-4o CoT Zero-shot 53.6 53.6 24.3 38.6 18.8 22.7 25.9 33.9
LLM-based dense retrievers
GritLM-7B SFT 24.8 32.3 18.9 19.8 17.1 13.6 17.8 20.6
GTE-QWEN-7B SFT 30.6 36.4 17.8 24.6 13.2 22.2 14.8 22.8
ReasonIR-8B SFT 26.2 31.4 23.3 30.0 18.0 23.9 20.5 24.8
Rerankers on BM25 Top-100 docs
RankGPT4 Zeroshot 33.8 34.2 16.7 27.0 22.3 27.7 11.1 24.7
RankZephyr-7b GPT4-distill 21.9 23.7 14.4 10.3 7.6 13.7 16.6 15.5
Rank-R1-14B GRPO (RL) 31.2 38.5 21.2 26.4 22.6 18.9 27.5 26.6
Rerankers on BM25+GPT-4o CoT Top-100 docs
Rank1-14B R1-distill 49.3 37.7 22.6 35.2 22.5 20.8 33.6 31.7
Rank-K-32B* R1-distill 50.8 49.4 28.2 46.0 27.3 30.5 31.9 37.9
Query expansion with BM25
HyDE-14B Zero-shot 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
LameR-14B Zero-shot 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
ThinkQE-14B (Ours) Zero-shot 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
Table 4: Results on the StackExchange domain of the BRIGHT benchmark in terms of nDCG@10. The best and the second best results across all models are in bold and underlined font, respectively. BM25+GPT-4o-CoT refers to applying BM25 to queries rewritten by GPT-4o with CoT reasoning traces included. All models except BM25+GPT-4o-CoT are performed on the original query. *Rank-K-32B performs computationally expensive listwise reranking over the top-20 documents.
Baselines.
[…]
On the BRIGHT benchmark, we consider three categories of baselines: (1) LLM-based embedding models such as GritLM-7B (Muennighoff et al., 2025), GTE-Qwen-7B (Li et al., 2023), and ReasonIR-8B Shao et al. (2025), all trained on massive amounts of retrieval data; (2) LLM-based rerankers, including RankGPT4 (zero-shot) (Sun et al., 2023), RankZephyr-7B (distilled from GPT-4) (Pradeep et al., 2023), Rank1-14B (distilled from DeepSeek-R1-685B) (Weller et al., 2025), Rank-R1-14B (trained via reinforcement learning) (Zhuang et al., 2025), and Rank-K-32B (distilled from DeepSeek-R1-685B) Yang et al. (2025). Rank1-14B, Rank-R1-14B, and Rank-K-32B explicitly incorporate a thinking process during reranking; and (3) Query expansion methods such as HyDE and LameR, use the same underlying model as ThinkQE but do not incorporate any explicit thinking process. [^3] We provide a detailed analysis of the no-thinking setting for fair comparison with ThinkQE in Section 4.1. Our method ThinkQE is evaluated in a zero-shot configuration across all datasets.
3.2 Main Results
[…]
On the BRIGHT benchmark, ThinkQE remains the strongest among zero-shot query expansion methods, achieving an average nDCG@10 of 34.9. While Rank-K-32B achieves the highest overall score (37.9), it relies on R1 distillation and listwise reranking over GPT-4o-augmented results, making it significantly more resource-intensive. In contrast, ThinkQE operates in a fully training-free setting and still outperforms several more expensive rerankers, including RankGPT4 (24.7) and Rank1-14B (31.7). Beyond its efficiency, ThinkQE delivers consistently strong performance across domains, ranking first or second in four sub-domains.
4 Analysis
In this section, we conduct a detailed analysis of ThinkQE on the StackExchange domain of the BRIGHT benchmark.
4.1 Impact of the Thinking Process
To evaluate the impact of the thinking process, we conduct two ablation studies on ThinkQE: (1) replacing the used model with its base version, QWEN-14B-Base, which has not been trained to produce reasoning traces, and (2) applying the NoThinking (Ma et al., 2025) method, where we prefill the response with a fabricated thinking block (i.e., <think>Okay, I think I have finished thinking.</think>) and allow the model to generate the answer directly from that point. As shown in Table 5, ThinkQE with thinking significantly outperforms both variants, underscoring the importance of generating thinking output. We use the NoThinking variant as the main baseline.
Model BRIGHT Avg.
QWEN-14B 27.6
QWEN-R1-14B w/o. thinking 29.8
QWEN-R1-14B w. thinking 32.5
Table 5: Impact on the thinking process. We compare three configurations: the base model QWEN-14B without any thinking involved, a NoThinking variant that bypasses actual thinking by prefilling a dummy thinking trace, and the ThinkQE model with the thinking process enabled.
4.2 Impact of Corpus Interaction
To evaluate the corpus interaction process, we compare ThinkQE to a baseline that performs all LLM expansions in a single round – referred to as parallel scaling. In contrast, ThinkQE uses corpus-interaction scaling, distributing expansions across multiple rounds with retrieval feedback. As shown in Figure 1, this interaction strategy consistently outperforms the static baseline, indicating that iterative refinement with evolving context is more effective than isolated expansions.
Figure 1: Impact of evolving corpus interaction process.
4.3 Impact on Expansion Accumulation and Redundancy Filter Mechanisms
We conduct a final ablation study on the two core components of the interaction process in ThinkQE: expansion accumulation, where query expansions from different rounds are concatenated to form the new query, and the semantic filter, which excludes top-retrieved documents from the previous round to encourage the introduction of new information. As shown in Table 6, both components are essential for maximizing performance. Disabling either mechanism leads to a noticeable performance drop, highlighting their complementary roles in refining the query and diversifying retrieved evidence across rounds.
Accum. Filter BRIGHT Avg.
34.2
33.4
34.9
Table 6: Impact of the expansion accumulation and redundancy filtering mechanisms.
5 Conclusion
We presented ThinkQE, a test-time query expansion method enhancing exploration and diversity through a thinking-based expansion process and evolving interactions with the corpus. Without requiring any training, ThinkQE consistently improves retrieval performance across multiple benchmarks by encouraging deeper coverage and adaptive refinement, offering a lightweight yet effective alternative to training-based dense retrievers and rerankers.
Limitations
The thinking process and evolving interaction process introduce higher inference-time latency and computational cost compared to single-shot expansion methods, which may limit its practicality in latency-sensitive or large-scale deployment scenarios. Furthermore, since our experiments focus exclusively on English web search tasks, the effectiveness of ThinkQE in multilingual settings remains unexplored.
Acknowledgements
This research was supported by project VI.Vidi.223.166 of the NWO Talent Programme which is (partly) financed by the Dutch Research Council NWO). We acknowledge the Dutch Research Council for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through project number NWO-2024.050.
Appendix A Appendix
A.1 Dataset Statistics
Details about the retrieval datasets are shown in Table 7.
Dataset #Test #Corpus
DL19 43 8,841,823
DL20 50 8,841,823
Biology 103 57,359
Earth Science 116 121,249
Economics 103 50,220
Psychology 101 52,835
Robotics 101 61,961
Stack Overflow 117 107,081
Sustainable Living 108 60,792
Table 7: Dataset Statistics
A.2 Detailed Results on the Impact of the Thinking Process
The detailed results across all domains on the impact of the thinking process are provided in Table 8.
Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
QWEN-BASE-14B 36.7 45.1 21.9 27.7 16.8 23.3 21.7 27.6
QWEN-R1-14B w/o. thinking 39.1 45.6 25.0 30.0 18.0 26.5 24.4 29.8
QWEN-R1-14B w. thinking 42.6 50.6 26.2 35.8 18.8 28.4 25.1 32.5
Table 8: Detailed results on the impact of the thinking process.
A.3 Detailed Results on the Impact of Evolving Corpus Interaction
The detailed results across all domains on the impact of the evolving corpus interaction are provided in Table 9.
#LLM calls Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
Parallel scaling
1 42.6 47.3 25.1 30.3 18.1 24.8 25.2 30.5
2 42.6 50.6 26.2 35.8 18.8 28.4 25.1 32.5
3 44.2 50.4 26.6 33.6 18.0 26.5 26.5 32.3
4 42.4 49.8 27.7 35.5 17.8 28.0 27.4 32.7
5 41.7 50.7 26.7 35.2 19.3 27.5 27.4 32.6
6 45.3 50.3 26.4 34.5 19.0 28.2 28.0 33.1
Corpus-interaction scaling
2 42.6 50.6 26.2 35.8 18.8 28.4 25.1 32.5
4 45.9 52.6 28.3 39.0 18.7 28.5 28.0 32.4
6 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
Table 9: Detailed results on the impact of the evolving corpus interaction.
A.4 Detailed Results on the Core Components of the Evolving Interaction Process
The detailed results across all domains on the impact of the expansion accumulation and redundancy filter mechanisms are provided in Table 10.
Accum. Filter Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
46.4 51.5 27.8 39.5 17.9 28.2 28.0 34.2
47.5 50.7 27.9 34.8 17.7 26.5 28.4 33.4
47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
Table 10: Detailed results on the impact of the expansion accumulation and redundancy filter mechanisms.

### v2

@@
 However, existing LLM-based methods often pay limited attention to exploration and result diversity.
-As illustrated in Table 1, we observe that current approaches – such as HyDE – tend to generate overly confident expansions that focus narrowly on a single interpretation of the input query.
+As illustrated in Table 1, we observe that current approaches, such as HyDE, tend to generate overly confident expansions that focus narrowly on a single interpretation of the input query.
 This behavior can be attributed to the model’s reliance on its internal knowledge and high-probability completions Yona et al.
@@
 We introduce ThinkQE, a query expansion framework that tightly integrates LLM-based thinking process with evolving corpus interaction.
+ThinkQE follows prior work in generating query expansions using retrieved documents but distinguishes itself through its design of thinking augmentation and iterative corpus interactions.
 The method is designed to enable exploration of the query space through thinking processes and evolving refinement based on retrieval feedback from the corpus.
@@
 At each round, an LLM performs thinking-augmented expansion based on the original query and newly retrieved documents from the corpus, which in turn informs subsequent retrieval and expansion steps.
-This section describes each component of the method in detail.
+The following subsections describe each component of the method in detail.
 2.1 Retrieving Initial Evidence from Corpus
@@
 ThinkQE Prompt
-Given a question "{$q$}" and its possible answering passages (most of these passages are wrong) enumerated as:
+Given a question "{qq}" and its possible answering passages (most of these passages are wrong) enumerated as:
 1. $\{d_{1}\}$; 2. $\{d_{2}\}$; 3. $\{d_{3}\}$
 please write a correct answering passage. Use your own knowledge, not just the example passages!
-[…]
+Table 2: Prompt used in ThinkQE for the thinking-based expansion process. $\{\cdot\}$ denotes the placeholder for the corresponding query and top-K documents.
 2.3 Evolution via Corpus Feedback
@@
 2.
-Redundancy Filtering: To promote diversity and avoid repetition, we exclude documents that (a) appear in the blacklist $\mathcal{B}_{t}$, or (b) were among the top-$K$ results in the previous round $\mathcal{D}_{t-1}$.
-We then select the top-$K$ documents from the remaining candidates: $\mathcal{D}_{t}^{\text{new}}=\text{TopK}(\mathcal{R}_{t}\setminus(\mathcal{B}_{t}\cup\mathcal{D}_{t-1})).$ The blacklist is updated to include all documents that were filtered out in this round.
+Redundancy Filtering: To promote diversity and avoid repetition, we exclude documents that (a) appear in the blocklist $\mathcal{B}_{t}$, or (b) were among the top-$K$ results in the previous round $\mathcal{D}_{t-1}$.
+We then select the top-$K$ documents from the remaining candidates: $\mathcal{D}_{t}^{\text{new}}=\text{TopK}(\mathcal{R}_{t}\setminus(\mathcal{B}_{t}\cup\mathcal{D}_{t-1})).$ The blocklist is updated to include all documents that were filtered out in this round.
 3.
@@
 This loop can be repeated for any number of rounds $T$, depending on resource constraints or desired depth.
-[…]
+Notably, as the query grows longer, successive expansions may dilute or replace the original intent.
+To mitigate this, we follow Zhang et al.
+(2024) and repeat the original query $n$ times in the final reformulation, with $n=\frac{\text{len}(\text{expansions})}{\text{len}(q_{0})\times\lambda},\quad\lambda=3.$ Here, $\text{len}(\text{expansions})$ refers to the total word count of all expansion segments, and $\text{len}(q_{0})$ is the word count of the original query.
+This repetition reinforces the core semantics of the original query during iterative refinement.
 Remark.
@@
 We set the total number of interaction rounds to 3, for a balance between efficiency and effectiveness.
+Besides the QWEN-R1-Distill-14B model, we also evalute ThinkQE on a variety of reasoning models, including QWEN3-8B, QWEN3-14B Yang et al.
+(2025a), OpenThinker2-7B Meta (2023), and Phi-4-Reasoning-14B Abdin et al.
+(2025).
 DL19 DL20
@@
 LameR 42.8 64.9 84.2 - - -
-ThinkQE-14B 45.9 68.8 89.3 43.9 64.7 87.8
+CSQE 43.6 63.4 87.6 - - -
+ThinkQE (ours)
+w. R1-14B 45.9 68.8 89.3 43.9 64.7 87.8
+w. QWEN3-8B 44.5 65.0 87.9 41.9 62.8 88.0
+w. QWEN3-14B 45.2 64.9 88.4 42.4 63.5 88.4
+w. OpenThinker2-7B 44.8 65.3 87.3 43.2 63.5 87.9
+w. Phi4-Reasoning-14B 44.0 65.0 87.1 43.0 63.9 87.2
 […]
@@
 Query expansion with BM25
-HyDE-14B Zero-shot 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
-LameR-14B Zero-shot 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
-ThinkQE-14B (Ours) Zero-shot 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
+HyDE-R1-14B Zero-shot 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
+LameR-R1-14B Zero-shot 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
+ThinkQE (ours)
+   R1-14B Zero-shot 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
+   QWEN3-8B Zero-shot 49.8 55.3 27.6 36.7 19.9 29.0 28.3 35.2
+   QWEN3-14B Zero-shot 51.5 53.2 27.8 37.2 22.0 16.1 27.5 33.6
+   OpenThinker2-7B Zero-shot 50.5 54.1 25.8 36.7 18.1 28.2 28.9 34.6
+   Phi-4-Reasoning-14B Zero-shot 51.8 53.5 29.7 38.5 21.8 29.3 27.7 36.0
 Table 4: Results on the StackExchange domain of the BRIGHT benchmark in terms of nDCG@10.
 The best and the second best results across all models are in bold and underlined font, respectively.
-BM25+GPT-4o-CoT refers to applying BM25 to queries rewritten by GPT-4o with CoT reasoning traces included.
-All models except BM25+GPT-4o-CoT are performed on the original query.
+All models are performed on the original query.
+BM25+GPT-4o-CoT refers to using BM25 retrieval results on queries rewritten by GPT-4o with chain-of-thought reasoning traces for reranking.
 *Rank-K-32B performs computationally expensive listwise reranking over the top-20 documents.
@@
 (2025), all trained on massive amounts of retrieval data; (2) LLM-based rerankers, including RankGPT4 (zero-shot) (Sun et al., 2023), RankZephyr-7B (distilled from GPT-4) (Pradeep et al., 2023), Rank1-14B (distilled from DeepSeek-R1-685B) (Weller et al., 2025), Rank-R1-14B (trained via reinforcement learning) (Zhuang et al., 2025), and Rank-K-32B (distilled from DeepSeek-R1-685B) Yang et al.
-(2025).
-Rank1-14B, Rank-R1-14B, and Rank-K-32B explicitly incorporate a thinking process during reranking; and (3) Query expansion methods such as HyDE and LameR, use the same underlying model as ThinkQE but do not incorporate any explicit thinking process.
+(2025b).
+Rank1-14B, Rank-R1-14B, and Rank-K-32B explicitly incorporate a thinking process during reranking; and (3) Query expansion methods such as HyDE and LameR, which use the same underlying model as ThinkQE but do not incorporate any explicit thinking process.
 [^3] We provide a detailed analysis of the no-thinking setting for fair comparison with ThinkQE in Section 4.1.
@@
 […]
-On the BRIGHT benchmark, ThinkQE remains the strongest among zero-shot query expansion methods, achieving an average nDCG@10 of 34.9.
-While Rank-K-32B achieves the highest overall score (37.9), it relies on R1 distillation and listwise reranking over GPT-4o-augmented results, making it significantly more resource-intensive.
+On the BRIGHT benchmark, ThinkQE remains the strongest among zero-shot query expansion methods, achieving an average nDCG@10 of 36.0 using the Phi-4-Reasoning-14B model.
+While Rank-K-32B achieves the highest overall score (37.9), it relies on R1 distillation and listwise reranking over GPT-4o-augmented retrieval results, making it significantly more resource-intensive.
 In contrast, ThinkQE operates in a fully training-free setting and still outperforms several more expensive rerankers, including RankGPT4 (24.7) and Rank1-14B (31.7).
@@
 Furthermore, since our experiments focus exclusively on English web search tasks, the effectiveness of ThinkQE in multilingual settings remains unexplored.
-Acknowledgements
+Acknowledgments
 This research was supported by project VI.Vidi.223.166 of the NWO Talent Programme which is (partly) financed by the Dutch Research Council NWO).
@@
 Sustainable Living 108 60,792
-Table 7: Dataset Statistics
-A.2 Detailed Results on the Impact of the Thinking Process
-The detailed results across all domains on the impact of the thinking process are provided in Table 8.
+Table 7: Dataset Statistics.
+A.2 Impact of the Thinking Process
+Results across all domains on the impact of the thinking process are provided in Table 10.
+A.3 Impact of Evolving Corpus Interaction
+Results across all domains on the impact of the evolving corpus interaction are provided in Table 11.
+A.4 Core Components of the Evolving Interaction Process
+Results across all domains on the impact of the expansion accumulation and redundancy filter mechanisms are provided in Table 12.
+A.5 Impact of $\lambda$
+Table 13 presents detailed results analyzing the impact of $\lambda$, which influences the repetition frequency of the original query during expansion.
+The results demonstrate that performance differences are small when varying $\lambda$ from 3 to 6.
+However, lower $\lambda$ values tend to cause excessive repetition of the original query, which generally hurts performance.
+A.6 Effectiveness-Efficiency Analysis on ThinkQE
+We provide latency and performance trade-off results comparing model size scaling, the thinking process, and multiple rounds in Table 14, using the DeepSeek-R1-Distill-Qwen-14B and DeepSeek-R1-Distill-Qwen-32B models evaluated on a single H100 GPU.
+Our results show that involving the thinking process and multiple rounds increases latency.
+However, scaling model size alone yields limited improvements relative to latency increase: moving from R1-14B without thinking (3.71 s/query, 29.8) to R1-32B without thinking (7.88 s/query) improves Bright Avg.
+by just +0.6 points (30.4) while more than doubling latency.
+Adding the thinking process, although it increases latency more, is substantially more effective: applying 1 round of thinking to R1-14B boosts performance by +2.7 points (32.5) compared to R1-14B without thinking, with latency rising to 15.40 s/query.
+Even with thinking enabled, scaling from R1-14B to R1-32B brings only a small additional gain (+0.4 points).
+Meanwhile, multi-round corpus interaction offers a more efficient path to higher effectiveness, with R1-14B 3-round reaching 34.9 (+2.4 over its 1-round version) and outperforming R1-32B 1-round.
+A.7 ThinkQE on Non-Web Search Datasets
+We evaluate ThinkQE on two additional non-web search datasets, TREC-Covid and Scifact.
+The results in Table 8 show that ThinkQE remains highly competitive and often outperforms the baselines.
+Note that both MILL and LameR are based on the powerful, closed-source GPT-3.5-Turbo model.
+Method TREC-Covid Scifact
+BM25 59.5 67.9
+Contriever (FT) 59.6 67.7
+HyDE 59.3 69.1
+Query2Doc 72.2 68.6
+MILL 75.3 71.4
+LameR 75.8 73.5
+CSQE 74.2 69.6
+ThinkQE 76.1 73.3
+Table 8: Results (NDCG@10) on non-web search datasets.
+A.8 Significance Test Results of ThinkQE
+We conduct significance testing by comparing ThinkQE with the two most relevant baselines we reimplemented on BRIGHT: HyDE and LameR using the same QWEN-R1-Distill-14B model.
+The results are presented in Table 15.
+Significance tests were performed using a t-test with a p-value threshold of 0.05.
+The results show that ThinkQE is significantly better than HyDE and LameR on 6 out of 7 and 5 out of 7 domains, respectively, demonstrating its effectiveness.
+A.9 Impact of Number of Rounds on Domain-Specific Dataset
+To further analyze the potential topic drift issue during the iterative process, we examine performance changes across one to three rounds using the domain-specific TREC-Covid dataset.
+The results, presented in Table 9, demonstrate that increasing from one to two rounds improves the performance, while extending to three rounds largely maintains performance.
+Round NDCG@10
+1 75.2
+2 76.2
+3 76.1
+Table 9: Performance across iterative rounds on the domain-specific TREC-COVID dataset.
 Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
@@
 Table 10: Detailed results on the impact of the thinking process.
-A.3 Detailed Results on the Impact of Evolving Corpus Interaction
-The detailed results across all domains on the impact of the evolving corpus interaction are provided in Table 9.
 #LLM calls Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
@@
 2 42.6 50.6 26.2 35.8 18.8 28.4 25.1 32.5
-4 45.9 52.6 28.3 39.0 18.7 28.5 28.0 32.4
+4 45.9 52.6 28.3 39.0 18.7 28.5 28.0 34.4
 6 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
 Table 11: Detailed results on the impact of the evolving corpus interaction.
-A.4 Detailed Results on the Core Components of the Evolving Interaction Process
-The detailed results across all domains on the impact of the expansion accumulation and redundancy filter mechanisms are provided in Table 10.
 Accum. Filter Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
@@
 47.3 52.5 29.2 40.0 19.3 28.0 27.9 34.9
-Table 10: Detailed results on the impact of the expansion accumulation and redundancy filter mechanisms.
+Table 12: Detailed results on the impact of the expansion accumulation and redundancy filter mechanism.
+$\mathbf{\lambda}$ Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
+1 38.2 47.5 25.0 32.1 18.4 25.5 24.0 30.1
+2 44.9 50.1 28.2 37.0 19.3 27.0 26.8 33.3
+3 46.7 51.2 29.4 39.3 20.8 28.3 28.4 34.9
+4 48.9 52.3 28.2 38.4 20.2 28.1 29.4 35.1
+5 49.7 52.5 29.5 37.7 19.6 28.9 27.6 35.1
+6 49.1 51.8 29.5 39.1 19.4 29.0 28.8 35.2
+Table 13: Detailed results on the impact of $\lambda$.
+Model Thinking Round Latency (second/query) Bright Avg.
+R1-14B No 1 3.71 29.8
+R1-32B No 1 7.88 30.4
+R1-14B Yes 1 15.40 32.5
+R1-32B Yes 1 30.53 32.9
+R1-14B Yes 3 45.44 34.9
+Table 14: Effectiveness-Efficiency Analysis on ThinkQE.
+Method Bio. Earth. Econ. Psy. Rob. Stack. Sus. Avg.
+HyDE 33.3 44.9 21.1 29.8 16.3 24.1 21.0 27.2
+LameR 35.1 46.1 23.7 31.0 17.7 26.4 25.3 29.3
+ThinkQE $47.3^{\dagger\ddagger}$ $52.5^{\dagger\ddagger}$ $29.2^{\dagger\ddagger}$ $40.0^{\dagger\ddagger}$ 19.3 $28.0^{\dagger}$ $27.9^{\dagger\ddagger}$ 34.9
+Table 15: Significance testing results.
+$\mathbf{\dagger}$ and $\mathbf{\ddagger}$ mean ThinkQE performs significantly better than HyDE and LameR, respectively, as determined by a t-test with p-value 0.05 as threshold.